State and prove Baire' s category theorem.
1
Expert's answer
2021-08-30T17:34:37-0400

Definition 1. A topological space X is said to be a Baire space, if for any given countable collection "A_{n}" of closed sets with empty interior in X, their union "\bigcup\limits_{n=1}^{\infty} A_n" also has empty interior in X.

In terms of open sets this property can be reformulated as follows:

Definition 2. A topological space X is said to be a Baire space, if for any given countable collection "G_{n}" of open dense subsets of X, their intersection "\bigcap\limits_{n=1}^{\infty} G_n" is dense.

Baire' s category theorem. Every complete metric space is a Bair space.

Proof. The set "\bigcap\limits_{n=1}^{\infty} G_n" is dense iff for every open set U, "\bigcap\limits_{n=1}^{\infty} G_n\cap U\ne \empty". We will construct a monotone sequence of closed balls "\bar B(x_{k+1},r_{k+1})\subset \bar B(x_{k},r_{k})" such that "\bar B(x_{k},r_{k})\subset \bigcap\limits_{n=1}^{k} G_n\cap U" and "r_k<2^{-k}". Then the sequence "\{x_n\}" will be a Cauchy sequence. Indeed, for all "\varepsilon>0" let "N\in\mathbb{N}" be arbitrary such that "2^{-N}<\varepsilon/2" and "n,m>N". Since "x_n\in \bar B(x_{n},r_{n})\in \bar B(x_{N},r_{N})" and "x_m\in \bar B(x_{m},r_{m})\in \bar B(x_{N},r_{N})" then "dist(x_n,x_m)<2r_N<2\cdot 2^{-N}<\varepsilon". Therefore, the sequence "\{x_n\}" is a Cauchy sequence and has a limit "x\in\bigcap\limits_{n=1}^{\infty}\bar B(x_{n},r_{n})\subset \bigcap\limits_{n=1}^{\infty} G_n\cap U". Therefore, "\bigcap\limits_{n=1}^{\infty} G_n\cap U\ne \empty".

A monotone sequence of closed balls "\bar B(x_{k+1},r_{k+1})\subset B(x_{k},r_{k})" can be constructed by induction, since for every "k\in\mathbb{N}" the set "G_{k+1}" is dense and open, and hence, has a non-empty intersection with non-empty open set "B(x_{k},r_{k})". Therefore, it contains some closed ball "\bar B(x_{k+1},r_{k+1})\subset \bar B(x_{k},r_{k})".

The proof is completed.